What is gravitational accelearation.

1
Expert's answer
2021-12-27T15:47:46-0500

Gravitational acceleration "g" is the rate of change of speed of a body with mass "m" released in the gravitational field of another object with mass "M". The gravitational acceleration can be found by Newton's law of universal gravitation in terms of distance between the centers of mass of the objects "R", mass of the object that creates the gravitational field, and universal gravitational constant "G":


"mg=G\dfrac{Mm}{R^2},\\\space\\
g=\dfrac{GM}{R^2}."